# Evaluating the Usability of Microgestures for Text Editing Tasks in Virtual Reality

### v1

@@ V-D Post-studies Questionnaires
We provided a post-studies questionnaire to evaluate participants’ (1) Perceived Fatigue, evaluated by using Borg Rating of Perceived Exertion (RPE) Scale (or Borg 6–20) [48], where 6 means “no exertion at all (rest)” and 20 represents maximal exertion, meaning the person is pushing themselves to their absolute physical limit; (2) Ease of Use; (3) Presence; and (4) Overall Preference, regarding the use of these two systems in both studies. The last three metrics were assessed via a single 7-point Likert scale, where 1 indicated “not at all” and 7 indicated “very much”, according to Gugenheimer et al. [49].
@@ VI-A2 Procedure
Participants in User Study 3 engaged in an open-ended scenario focused on information gathering through web browsing and note-taking, without specific tasks or wrong actions. The study was conducted in a virtual environment, where participants viewed two screens simultaneously: the left screen displayed a web browser with information about various travel-related topics (e.g., mountain locations, opening hours, altitudes, etc.), and the right screen showed a notes application with sample questions (e.g., what is the altitude of the mountain). Participants were required to find the relevant answers from the web browser and then add the proper information into the notes application.
@@ VI-B1 Qualitative Feedback
The Pragmatic Quality scores showed no significant difference between the Baseline ($M=1.25$, $SD=1.31$) and microGEXT ($M=1.9$, $SD=0.59$; $U=39.50$, $p=0.4471$, $r=0.210$). However, significant differences were observed for Hedonic Quality, where microGEXT ($M=1.9$, $SD=0.60$) outperformed the Baseline ($M=-0.425$, $SD=1.31$; $U=9.00$, $p=0.0021$, $r=0.820$). Similarly, Overall User Experience scores were significantly higher for microGEXT ($M=1.9$, $SD=0.51$) compared to the Baseline ($M=0.4125$, $SD=1.09$; $U=14.50$, $p=0.0080$, $r=0.710$).
@@
Hedonic Quality -0.425 ± 1.31 (↓) 1.9 ± 0.60 (**) (↑)
Overall Experience 0.4125 ± 1.09 (↓) 1.9 ± 0.51 (**) (↑)
@@ VII Discussion, Limitations, and Future Work
A key finding across all studies is the significant reduction in physical demand and fatigue when using microGEXT. Traditional VR input methods, especially those that rely on large arm movements, are prone to causing users fatigue over time, as exemplified by the gorilla arm syndrome [3, 4]. microGEXT addresses this challenge effectively through small, precise hand movements that reduce the physical strain typically associated with text editing tasks in VR. The NASA-TLX results from Studies 1 and 2 support this, showing that microGEXT consistently lowered physical and mental demand, while also reducing frustration, particularly in tasks requiring high precision like text selection and deletion.
However, it is also clear from the results that microGEXT has its limitations, especially for complex tasks requiring long text selections. For example, in Study 2, we found that microGEXT was significantly slower than the Baseline system for paragraph and two-paragraph text selections. This suggests that while microGEXT excels at handling short, quick interactions, its two-step interaction process (wrist rotation to change modes followed by gesture execution) can introduce delays for more complex operations. This highlights the importance of balancing simplicity and efficiency in gesture design. Future iterations of the system could explore optimizing the interaction flow for long-range selections, possibly through more dynamic or context-sensitive gesture controls.
While the accuracy of the microGEXT recognition system was effective in the studies, we see potential to further improve its performance. The recognition model was trained and calibrated using 12 datasets from 6 subjects, each performing the gestures twice. While a larger and more diverse dataset could improve the model’s ability to generalize to new users, the addition of new gestures would require further data collection and model retraining, which poses a scalability challenge. Future work could also explore the use of automated data augmentation techniques or transfer learning methods to improve recognition accuracy without requiring a vast amount of additional training data.
The current microGEXT system incorporates eight dynamic gestures designed for common text editing tasks. As the number of gestures in the system increases, relying solely on gesture-based interaction becomes increasingly challenging. A larger gesture set makes it more difficult for users to remember and recognize gestures, while also requiring more training data, thus complicating usability and increasing the demand for system resources. To address this, we see interesting future research in investigating how to build systems that could employ a hybrid interaction model that combines microgestures with other input methods, such as voice commands or contextual menus, allowing users to switch between interaction modes depending on task complexity.
Finally, it would be interesting to study the adoption of microGEXT in deployment studies, which would allow a nuanced understanding of barriers, appropriation moves, and other aspects that emerge through interaction with technologies in everyday life [50].
@@ VIII Conclusion
This paper introduced microGEXT, a microgesture-based system for text editing in virtual reality (VR), designed to address challenges like user fatigue, precision limitations, and social accessibility. We evaluated its usability, efficiency, and user experience through three user studies.
Study 1 ($N=20$) compared microGEXT with a Baseline gesture-plus-menu system, showing significant reductions in edit time for commands like Cut, Delete, and Select All, along with improved user experience. Study 2 ($N=20$) focused on text range selection, revealing better performance for shorter tasks and reduced physical and mental demand for longer ones. Study 3 ($N=10$) highlighted microGEXT’s effectiveness in open-ended tasks, minimizing fatigue and supporting seamless task transitions.
Overall, microGEXT demonstrates how microgesture interaction has the potential to bestow VR input methods with interactions that reduce fatigue and increase user satisfaction while retaining all or most performance of traditional mid-air interfaces. We see promising future work in exploring how to further improve recognition accuracy, expand the gesture set, and study and optimize the system in prolonged usage scenarios in real-world applications. To support further research in this area, we provide our open-source code and anonymized datasets as supplemental materials.
@@ Acknowledgments
Xiang Li is supported by the China Scholarship Council (CSC) International Cambridge Scholarship (No. 202208320092). Per Ola Kristensson is supported by the EPSRC (grants EP/S027432/1 and EP/W02456X/1).

### v2

@@ V-D Post-studies Questionnaires
We provided a post-studies questionnaire to evaluate participants’ (1) Perceived Fatigue, evaluated by using Borg Rating of Perceived Exertion (RPE) Scale (Borg 6–20) [52], where 6 means “no exertion at all (rest)” and 20 represents maximal exertion, meaning the person is pushing themselves to their absolute physical limit; (2) Ease of Use; (3) Presence; and (4) Overall Preference, regarding the use of these two systems in both studies. The last three metrics were assessed via a single 7-point Likert scale, where 1 indicated “not at all” and 7 indicated “very much.”
@@ VI-A2 Procedure
Participants in User Study 3 engaged in an open-ended scenario focused on information gathering through web browsing and note-taking, without specific tasks or wrong actions. The study was conducted in a virtual environment, where participants viewed two screens simultaneously: the left screen displayed a web browser with information about various travel-related topics (e.g., mountain locations, opening hours, altitudes, etc.), and the right screen showed a notes application with sample questions (e.g., what is the altitude of the mountain). Participants were required to find the relevant answers from the web browser and then add the proper information into the notes application. The study protocol was also approved by the Cambridge University Engineering Department Ethics Committee (#452).
@@
Figure 7: Screenshot from User Study 3, showing the scenario where participants performed information gathering and note-taking tasks. The left side of the screen displays a web browser with travel-related information about the Alps, and the right side shows a notes application with sample questions that participants needed to answer by extracting relevant data from the web page.
@@ VI-B1 Qualitative Feedback
The Pragmatic Quality scores showed no significant difference between the Baseline ($M=1.25$, $SD=1.31$) and microGEXT ($M=1.9$, $SD=0.59$; $U=39.50$, $p=0.4471$, $r=0.210$). However, significant differences were observed for Hedonic Quality, where microGEXT ($M=1.9$, $SD=0.60$) outperformed the Baseline (M=0.425M=-0.425, $SD=1.31$; $U=9.00$, $p=0.0021$, $r=0.820$). Similarly, Overall User Experience scores were significantly higher for microGEXT ($M=1.9$, $SD=0.51$) compared to the Baseline ($M=0.4125$, $SD=1.09$; $U=14.50$, $p=0.0080$, $r=0.710$).
@@
Hedonic Quality -0.43 ± 1.31 (↓) 1.9 ± 0.60 (**) (↑)
Overall Experience 0.41 ± 1.09 (↓) 1.9 ± 0.51 (**) (↑)
@@ VII Discussion, Limitations, and Future Work
A key finding across all studies is the significant reduction in physical demand and fatigue when using microGEXT. Traditional VR input methods, especially those that rely on large arm movements, are prone to causing users fatigue over time, as exemplified by the gorilla arm syndrome [3, 4]. microGEXT addresses this challenge effectively through small, precise hand movements that reduce the physical strain typically associated with text editing tasks in VR. The NASA-TLX results from Studies 1 and 2 support this, showing that microGEXT consistently lowered physical and mental demand, while also reducing frustration, particularly in tasks requiring high precision like text selection and deletion. In future work, fatigue evaluation could benefit from computational modeling approaches such as NICER [53] or adaptive interventions like AlphaPIG [54], which provide predictive and real-time measures of physical strain in XR.
Another dimension concerns gesture memorability. Prior elicitation studies [14] and large-scale surveys of gesture sets [17] have shown that as the number of gestures increases, users experience difficulties in recall and consistency. Recent work on transferable microgestures across contexts [16] reinforces the need to balance ergonomic feasibility with cognitive load. Our studies suggest that microGEXT’s compact set is memorable, but expanding beyond eight commands may require hybrid designs or mnemonic aids to prevent usability degradation.
While the recognition system achieved high accuracy, our dataset was limited in size and diversity, raising potential concerns about generalization to broader user populations. Adding new gestures would also necessitate retraining, limiting scalability. Future work could leverage data augmentation or transfer learning to expand coverage without exhaustive new data collection. Cross-population studies involving users with different cultural or ergonomic backgrounds would further strengthen generalizability.
The current system relies solely on microgestures, which highlights both strengths and limitations. While effective for short-range editing, performance slowed in longer-range selection tasks (Study 2). This suggests the need for hybrid modalities that combine microgestures with other inputs such as voice commands or contextual menus, allowing users to flexibly adapt to task complexity. Exploring such multimodal synergies could preserve efficiency while maintaining low fatigue.
Finally, we see important opportunities in deployment studies. Lab-based tasks capture controlled performance metrics but miss appropriation, adaptation, and long-term use patterns that emerge in everyday workflows. Inspired by recent work on extended deployment in XR [55], future research could examine microGEXT in authentic productivity contexts to uncover barriers, adoption strategies, and sustained benefits.
In summary, microGEXT highlights the potential of microgestures for efficient and low-fatigue VR text editing, while also surfacing broader design and methodological challenges around fatigue evaluation, memorability, dataset scalability, and multimodal integration.
@@ VIII Conclusion
This paper introduced microGEXT, a microgesture-based system for text editing in VR, designed to mitigate well-known challenges such as user fatigue, precision limitations, and social accessibility. We evaluated its usability, efficiency, and overall user experience through three controlled studies.
Study 1 ($N=20$) compared microGEXT with a commercially inspired Baseline gesture-plus-menu system, showing significant reductions in edit time for commands such as Cut, Delete, and Select All, alongside improved user experience. Study 2 ($N=20$) examined text range selection, where microGEXT performed comparably to the Baseline for short selections while reducing physical and mental demand in longer tasks. Study 3 ($N=10$) demonstrated that microGEXT was intuitive and fatigue-free in open-ended workflows, supporting seamless task switching during extended use.
Taken together, our findings demonstrate that microgestures can provide a lightweight yet powerful alternative to conventional VR text editing techniques. By reducing fatigue and enhancing user satisfaction while retaining most of the performance of mid-air interfaces, microGEXT highlights the potential of subtle input for everyday VR productivity. Future work should explore refining recognition accuracy, extending the gesture vocabulary, and investigating long-term adoption in real-world scenarios, including hybrid modalities that combine microgestures with complementary input channels. To facilitate further research, we release our open-source code and anonymized datasets as supplemental materials.
@@ Acknowledgments
Xiang Li is supported by the China Scholarship Council (CSC) International Cambridge Scholarship (No. 202208320092). Wei He is supported by the HKUST(GZ) Research Studentship. Per Ola Kristensson is supported by the EPSRC (grants EP/S027432/1 and EP/W02456X/1).